# Reward-Aware Proto-Representations in Reinforcement Learning

### v1

@@ B.4 Convergence Proof for Dynamic Programming Algorithm
Let $\mathcal{S}_{N}$ denote the set of non-terminal states. Assume $r(s)<0\ \forall s\in\mathcal{S}_{N}$. Let ${\mathbf{R}}=\operatorname{diag}(\exp(-{\mathbf{r}}/\lambda))$, where ${\mathbf{r}}$ is the vector of all state rewards. Let ${\mathbf{Z}}_{0}={\mathbf{R}}^{-1}$. The update rule
@@
B.5 Derivation of the MER
In this subsection, we provide the full derivation of the maximum entropy representation (MER), which is the proto-representation in the maximum entropy RL framework.
To derive the MER for deterministic transitions, we can simply formulate the entropy as $\mathcal{H}\big{(}\pi(\cdot|S_{t})\big{)}$. However, to derive the MER for stochastic transitions, we need to formulate the entropy slightly differently. We assume that at every time step $t$, the agent receives a reward of $\tilde{r}(S_{t})=r(S_{t})+\lambda\mathcal{H}\big{(}p^{\pi}(\cdot|S_{t})\big{)}$, where $p^{\pi}(\cdot|S_{t})$ denotes the transition probabilities over the successor states of $S_{t}$ given policy $\pi$, and $\lambda>0$ determines the relative importance of the entropy term.
@@ Appendix C More Experiment Details
We provide additional experiment details omitted from the main text due to space constraints.
@@ C.1 Numerical Considerations
Reward shaping and option discovery experiments involve the top eigenvector of the DR. We describe numerical considerations when performing eigendecomposition of the DR. First, in practice, we perform eigendecomposition of the symmetrized DR, $\operatorname{Sym}({\mathbf{Z}})=({\mathbf{Z}}+{\mathbf{Z}}^{\top})/2$, to ensure real eigenvalues and eigenvectors. Second, as the DR involves exponentiating negative rewards (see Eq. 7), the magnitude of the DR entries, especially the off-diagonal entries, can become very small. To mitigate the resulting numerical issues, we use the library python-flint. [^1] https://github.com/flintlib/python-flint, MIT license Note, however, that the improved precision comes at the cost of increased runtime.
@@ C.2 Reward Shaping
Figure 6 shows reward-shaping results in the variations of the environments shown in Figure 1 without low-reward regions. In these environments, we do not observe a significant performance difference between potential-based reward shaping using the DR and the SR. This is because in these environments, all states apart from the terminal state have the same reward. Although Theorem 3.1 cannot be directly applied since the terminal state has a different reward (00) than the rest of the states ($-1$), it is reasonable to expect that the DR and the SR have similar eigenvectors in this setting, leading to similar results.
@@ C.3 Option Discovery
We share the following hyperparameters for RACE and CEO: $\alpha_{0}=0.1$, $\gamma_{0}=0.99$, $N_{\text{step}}=100$, $N_{\text{iter}}=50$ for grid task and four rooms, and $N_{\text{iter}}=120$ for grid room and grid maze. We use $\gamma=0.99$ for learning the SR, $\lambda=1.3$ and the uniform random policy as the default policy for learning the DR, and sweep over the remaining hyperparameters, as described in Table 2. We perform 10 independent runs for each hyperparameter setting.
@@
              $\mathcal{D}_{\text{curr}}\longleftarrow\mathcal{D}_{\text{curr}}\|(s,a,r,s^{\prime})$triangleright Append transition to current dataset
@@
                  $\mathcal{D}_{\text{curr}}\longleftarrow\mathcal{D}_{\text{curr}}\|(s,a,r,s^{\prime})$triangleright Append transition to current dataset
@@
For RACE+Q and CEO+Q, we follow the same hyperparameter search procedure as before (see Table 2). For the Q-learning baseline, we perform a grid search over the Q-value initialization ($[-1000,-100,-10,0]$), $\epsilon$ for $\epsilon$-greedy exploration ($[0.01,0.05,0.1,0.15,0.2]$), and step size ($[0.01,0.03,0.1,0.3,1]$). We run 10 seeds for each hyperparameter setting, and after identifying the best hyperparameters, re-run 50 seeds to avoid maximization bias. As we do not see much performance difference in simpler environments, we introduce two new environments that are larger versions of Grid Room (Grid Room (L)) and Grid Maze (Grid Maze (L)), shown in Figure 10.
@@ C.4 Count-Based Exploration
To facilitate the learning and application of the DR for exploration in the Riverswim and Sixarms environments, we rescale the environment rewards to lie within the range $[-1,0]$. Importantly, this rescaling is applied solely for learning the DR; the Sarsa agent continues to operate using the original, unscaled rewards for action-value estimation. Note that we use the DR defined for state-action-dependent rewards, $\bar{{\mathbf{Z}}}$.
@@ C.5 Transfer
For the SR, we have to represent the reward function for all states. There are three types of states in the environment: empty state (indicated by white tile), low-reward state (indicated by red tile), and goal state (indicated by green tile). All empty states have the same reward of $-1$, and all low-reward states have the same reward of $-20$. Therefore, we represent state features as a 6-dimensional one-hot vector, where one entry is activated when the state is an empty state, one entry is activated when the state is a low-reward state, and the remaining four entries correspond to the four terminal states.
@@ Appendix D Statistical Significance
We report 95% confidence intervals or directly visualize all independent runs if possible. The 95% confidence intervals capture randomness across independent runs, where the primary source of randomness is random exploration, e.g., $\epsilon$-greedy exploration. In transfer learning experiments, randomness also arises from sampling terminal state reward configurations. As we perform a large number of independent runs in our experiments ($N\geq 50$), we assume the sampling distribution is normal and compute the 95% confidence interval as ±$\pm 1.96\frac{\hat{\sigma}}{\sqrt{N}}$, where $\hat{\sigma}$ is the sample standard deviation.

### v2

@@ B.4 Convergence Proof for Dynamic Programming Algorithm
[…]
@@
B.5 Proto-Representation in Maximum Entropy RL
Definition 4.1.
The MER differs from the DR (Equation 4) in that it uses the adjacency matrix rather than the default policy’s transition probabilities. This connection arises because maximum entropy RL and linearly solvable MDPs are closely related as one can be transformed into the other [11]. As a result, most of our contributions readily extend to the MER. Preliminary analyses showed that the MER also behaves similarly to the DR, so we focus our experiments on the DR. Still, the MER and DR are distinct proto-representations with different formulations, assumptions, and value functions, which can lead to different optimal policies. While we did not observe substantial differences in the settings we considered, exploring them in other regimes remains an open direction.
We now provide the derivation of the MER. To derive the MER for deterministic transitions, we can simply formulate the entropy as $\mathcal{H}\big(\pi(\cdot|S_{t})\big)$. However, to derive the MER for stochastic transitions, we need to formulate the entropy slightly differently. We assume that at every time step $t$, the agent receives a reward of $\tilde{r}(S_{t})=r(S_{t})+\lambda\mathcal{H}\big(p^{\pi}(\cdot|S_{t})\big)$, where $p^{\pi}(\cdot|S_{t})$ denotes the transition probabilities over the successor states of $S_{t}$ given policy $\pi$, and $\lambda>0$ determines the relative importance of the entropy term.
@@ Appendix C More Experiment Details
We provide additional experiment details omitted from the main text due to space constraints. [^1] The code for this paper can be found at: github.com/httse9/Reward-Aware-Proto-Representations
@@ C.1 Numerical Considerations
Reward shaping and option discovery experiments involve the top eigenvector of the DR. We describe numerical considerations when performing eigendecomposition of the DR. First, in practice, we perform eigendecomposition of the symmetrized DR, $\operatorname{Sym}({\mathbf{Z}})=({\mathbf{Z}}+{\mathbf{Z}}^{\top})/2$, to ensure real eigenvalues and eigenvectors. Second, as the DR involves exponentiating negative rewards (see Eq. 7), the magnitude of the DR entries, especially the off-diagonal entries, can become very small. To mitigate the resulting numerical issues, we use the library python-flint. [^2] https://github.com/flintlib/python-flint, MIT license Note, however, that the improved precision comes at the cost of increased runtime. As a larger $\lambda$ has the effect of reducing the magnitude of negative trajectory returns in the DR (see Eq. 7), it can alleviate numerical instabilities. In our experiments, we initially started with $\lambda=1$, and slowly increased $\lambda$ by 0.1 until we settled on $\lambda=1.3$, which, paired with the use of python-flint, resulted in no numerical issues.
@@ C.2 Reward Shaping
Figure 6 shows reward-shaping results in the variations of the environments shown in Figure 1 without low-reward regions. In these environments, we do not observe a significant performance difference between potential-based reward shaping using the DR and the SR. This is because in these environments, all states apart from the terminal state have the same reward. Although Theorem 3.1 cannot be directly applied since the terminal state has a different reward (0) than the rest of the states (1-1), it is reasonable to expect that the DR and the SR have similar eigenvectors in this setting, leading to similar results.
@@ C.3 Option Discovery
We share the following hyperparameters for RACE and CEO: $\alpha_{0}=0.1$, $\gamma_{0}=0.99$, $N_{\text{step}}=100$, $N_{\text{iter}}=50$ for grid task and four rooms, and $N_{\text{iter}}=120$ for grid room and grid maze. We use $\gamma=0.99$ for learning the SR, $\lambda=1.3$ for learning the DR, and sweep over the remaining hyperparameters, as described in Table 2. We perform 10 independent runs for each hyperparameter setting.
@@
     $\mathcal{D}_{\text{curr}}\longleftarrow\mathcal{D}_{\text{curr}}\|(s,a,r,s^{\prime})$ $\triangleright$ Append transition to current dataset
@@
      $\mathcal{D}_{\text{curr}}\longleftarrow\mathcal{D}_{\text{curr}}\|(s,a,r,s^{\prime})$ $\triangleright$ Append transition to current dataset
@@
For RACE+Q and CEO+Q, we follow the same hyperparameter search procedure as before (see Table 2). For the Q-learning baseline, we perform a grid search over the Q-value initialization ([1000,100,10,0][-1000,-100,-10,0]), $\epsilon$ for $\epsilon$-greedy exploration ($[0.01,0.05,0.1,0.15,0.2]$), and step size ($[0.01,0.03,0.1,0.3,1]$). We run 10 seeds for each hyperparameter setting, and after identifying the best hyperparameters, re-run 50 seeds to avoid maximization bias. As we do not see much performance difference in simpler environments, we introduce two new environments that are larger versions of Grid Room (Grid Room (L)) and Grid Maze (Grid Maze (L)), shown in Figure 10.
@@ C.4 Count-Based Exploration
To facilitate the learning and application of the DR for exploration in the Riverswim and Sixarms environments, we rescale the environment rewards to lie within the range [1,0][-1,0]. Importantly, this rescaling is applied solely for learning the DR; the Sarsa agent continues to operate using the original, unscaled rewards for action-value estimation. Note that we use the DR defined for state-action-dependent rewards, $\bar{{\mathbf{Z}}}$.
@@ C.5 Transfer
For the SR, we have to represent the reward function for all states. There are three types of states in the environment: empty state (indicated by white tile), low-reward state (indicated by red tile), and goal state (indicated by green tile). All empty states have the same reward of 1-1, and all low-reward states have the same reward of 20-20. Therefore, we represent state features as a 6-dimensional one-hot vector, where one entry is activated when the state is an empty state, one entry is activated when the state is a low-reward state, and the remaining four entries correspond to the four terminal states.
@@ Appendix D Statistical Significance
We report 95% confidence intervals or directly visualize all independent runs if possible. The 95% confidence intervals capture randomness across independent runs, where the primary source of randomness is random exploration, e.g., $\epsilon$-greedy exploration. In transfer learning experiments, randomness also arises from sampling terminal state reward configurations. As we perform a large number of independent runs in our experiments ($N\geq 50$), we assume the sampling distribution is normal and compute the 95% confidence interval as $\pm 1.96\frac{\hat{\sigma}}{\sqrt{N}}$, where $\hat{\sigma}$ is the sample standard deviation.